L^{A}TEX News

Issue 16, December 2003

Anniversary news

This anniversary Issue 16 takes a brief look into the future work of the L^{A}TEX3 Project Team, both short and and longer range. Please let us know if you want to get involved with us in any of this work (see below).

An overview of the 10th Anniversary Release, dated 2003/12/01, is can be found in Issue 15.

TLC2: The L^{A}TEX Companion – 2nd edition!

Since you are reading this newsletter, there is a good chance that you, or a friend, has already bought this encyclopedic volume: the incomparable Second Edition of this work that is every L^{A}TEXie’s ultimate lucky charm.

If by some chance you have not yet purchased your own copy then get into training, get shopping, and get flexing your muscles (both physical—it’s 1100+ pages, and intellectual) by using it to discover masses of invaluable ‘insider information’ about:

• the latest release of Standard L^{A}TEX;

• over 200 extension packages;

• plus related software and systems.

For more information on this all new (??… OK, not all, but over 90%!!), all accurate (we hope!) 10th Anniversary Edition, check out

http://www.awprofessional.com/titles/0201362996.

Future maintenance

We are currently exploring how best to support the very large and rapidly growing community of individuals, organisations and enterprises that depend on the robustness and availability of the current standard L^{A}TEX distribution. Although we remain firmly resolved not to make changes in the base distribution (the kernel) of Standard L^{A}TEX, there is still much that needs doing to maintain its reliability and utility and to keep up the necessary level of communication with users and supporters. Also, as with all advanced software systems, bugs are still turning up occasionally so some fixes are still essential.

One major impediment to providing adequate service levels in this area is, of course, the difficulties inherent in obtaining the time and commitment of skilled minds—hence the appeal above to anyone interested in getting involved.

LPPL certification

There are still some outstanding diplomatic tasks around the L^{A}TEX Project Public Licence: these include e.g., getting it ‘OSF certified’ and ensuring that it gains more support and wider use, even in the FSF world where it has long been tolerated.

Use of ε-TEX/pdfTEX

We expect that within the next two years, releases of L^{A}TEX will change modestly in order to run best under an extended TEX engine that contains the ε-TEX primitives, e.g., ε-TEX or pdfTEX. The details of this possible upgrade need further work so we are not making a definite announcement yet.

Although the current release does not require ε-TEX features, we certainly recommend using an extended TEX, especially if you need to debug macros.

End of ‘autoload’ support

As computer systems generally grow in capacity, requirements change and so we believe that the autoload variant of L^{A}TEX is no longer required. Thus, although the code remains it is no longer supported. We hope this does not cause any problems.

New models, new code

In the period 1999–2001 we published many results of our work over the previous decade on the development of new concepts and models for automated typesetting based on TEX as the underlying platform. These can be

found at http://www.latex-project.org/papers/ and
http://www.latex-project.org/code/experimental/.

Since then a very large proportion of the The Team’s efforts have been diverted to provide the core author team for TLC2, which provides over 1000 pages of carefully researched and tested documentation of many aspects of the vast world of L^{A}TEX related software that was developed over that same time period and that continues to grow and improve prodigiously.

Completion of that task … until TLC3!! … presents the possibility of getting back to this more exciting development work, or even to more radical work on non-TEX-based models and implementations.

Of course, any such ideas are predicated on our ability to organise (with you, we hope) an efficient but responsive maintenance and support system for Standard L^{A}TEX.

L^{A}TEX News, and the L^{A}TEX software, are brought to you by the L^{A}TEX3 Project Team; Copyright 2003, all rights reserved.